OpenStudy (anonymous):

given the equation y - 3/4 (x + 8) in point slop form, identify the equation of the same line in slope-intercept form.

OpenStudy (anonymous):

−3/4x + y = 10 3x − 4y = −40 y = 3/4x + 12 y = 3/4x + 10

OpenStudy (skullpatrol):

First of all this " y - 3/4 (x + 8)" is NOT an equation. Can you tell me why?

OpenStudy (skullpatrol):

I believe the answer is y = (3/4)x +10, given the equation y - 4 = (3/4)(x + 8).